# Meta Curvature-Aware Minimization for Domain Generalization

### v1

@@ Appendix D Additional Experiments
[…]
  Algorithm PACS VLCS OfficeHome TerraInc DomainNet Average
Adam + SWAD $86.8$ $79.1$ $70.1$ $46.5$ $44.1$ $65.3$
AdamW + SWAD $87.0$ $78.5$ $70.8$ $46.9$ $45.0$ $65.6$
SGD + SWAD $85.2$ $79.1$ $71.0$ $46.7$ $42.8$ $65.0$
FAD + SWAD $88.5$ 79.8 71.8 $47.5$ $45.0$ $66.5$
MeCAM (Ours) + SWAD 88.9 79.8 $70.8$ 48.9 45.9 66.9
@@
D.1 Extensibility Analysis
To evaluate the extensibility of our MeCAM, we integrated it with SWAD [9], which is a typical DG method based on ensemble learning, and repeated experiments on five DG datasets. The optimal hyperparameter configurations listed in Table B were directly applied to MeCAM. For SWAD, the hyperparameters were set to their default values: ‘n_converge’=3, ‘n_tolerance’=6, and ‘tolerance_ratio’=0.3. Therefore, no additional hyperparameter search was performed. The results are presented in Table XII. It shows that our MeCAM integrated with SWAD outperforms other combinations on four datasets and also achieves the best overall performance, demonstrating the superior extensibility of our MeCAM.
D.2 Ablation Study on Hyperparameters

### v2

@@ Appendix D Additional Experiments
Table XII: Ablation study on various combinations of optimization objectives in our MeCAM. Average accuracy and standard error ($mean_{\pm std}$) of our MeCAM and other variants are calculated across three trials on PACS, VLCS, and OfficeHome. The best results are highlighted in bold.
$85.0_{\pm 0.5}$ $78.2_{\pm 0.5}$ $67.8_{\pm 0.2}$ $77.0$
$15.4_{\pm 0.3}$ $47.0_{\pm 0.3}$ $1.6_{\pm 0.1}$ $21.3$
$46.1_{\pm 5.2}$ $44.2_{\pm 2.9}$ $2.3_{\pm 0.2}$ $30.9$
$86.4_{\pm 1.3}$ $\textbf{80.3}_{\pm 0.5}$ $69.7_{\pm 0.2}$ $78.8$
$86.7_{\pm 0.5}$ $79.7_{\pm 0.3}$ $69.5_{\pm 0.5}$ $78.6$
$87.4_{\pm 0.1}$ $79.4_{\pm 0.1}$ $69.8_{\pm 0.2}$ $78.9$
$\textbf{88.3}_{\pm 0.3}$ $\textbf{80.3}_{\pm 0.3}$ $\textbf{70.4}_{\pm 0.4}$ 79.7
@@
D.1 Ablation Study
To better understand the contributions of each component in MeCAM, we conducted an ablation study on the three terms of MeCAM’s optimization objective (see Eq. (11) in our manuscript). The results were displayed in Table XII. Please note that when only $\alpha(f(\theta+\delta)-f(\theta))$ or $\beta(m(\theta-\delta)-f(\theta))$ is used as the optimization objective, we set $\alpha$ or $\beta$ to 1. The results highlight the following insights: (1) using only $f(\theta)$ degenerates MeCAM to ERM, which does not effectively reduce sharpness or curvature; (2) relying solely on the surrogate gap of SAM or meta-learning results in significant performance degradation, particularly on the OfficeHome dataset, indicating that minimizing sharpness alone is insufficient; (3) combining the surrogate gaps of SAM or meta-learning with $f(\theta)$ alleviates this degradation and improves generalization performance; and (4) the best performance is achieved when all three terms are jointly used in the optimization objective, suggesting the effectiveness of MeCAM. Furthermore, it is worth noting that combining the surrogate gap of SAM with meta-learning not only resolves the degradation observed when using either of them individually, but also outperforms the use of $f(\theta)$ alone. We argue that the optimization objectives of these two terms are complementary, and their combination effectively mitigates limitations of each other.
D.2 Discussion on Hyperparameters